Grade 12Algebra

Solve the following equations:
0 = 16 + 4(m – 6)............

Pawan Prajapati
5 Years agoGrade 12

1 Answer

Harshit Singh
5 Years ago
Dear Student

By transposing 16 from RHS to LHS it becomes–16
= 0–16 = 4(m - 6)
= - 16 = 4(m - 6)
Let us divide both the side by 4,
= - 16/4 = (4(m - 6))/ 4
= - 4 =m - 6
By transposing - 6 from RHS to LHS it becomes 6
= - 4 +6 =m
= m =2

Thanks